Crew 263 Sol Summary Report 28-02-2022
Summary Title: Back to work
Author’s name: Cerise Cuny
Mission Status: Nominal
Sol Activity Summary: Surprisingly after a day of rest, we felt more tired this morning when we woke up than during the previous week. The two morning EVAs went smoothly. Meanwhile, the crew Engineer and I tested the broken earpiece with several radios and modes. We spotted some exposed wires which may be the source of the issue. As the earpiece has not been subject to any violent/unusual use, maybe it is very sensitive. To make sure we do not damage any other earpiece, we will think about ways to attach the wire so that it cannot get caught in anything while on EVA. Our astronomer worked on astrophotography images and tried to retrace the light curve of an exoplanet transit using the data from the New Mexico remote telescope.
We had lunch all together, as usual: we deeply value these moments when we can share the latest development in our experiments, our feelings about being on Mars, far from our family and friends, how confinement and isolation impacts our mood, etc.
In the afternoon, our GreenHab officer started a spiruline experiment, which reminded her of her role even with the GreenHab being closed. We prepared our EVA plans for tomorrow. Some of us did artistic activities (water color painting and knitting).
Anomalies in work: None.
Weather: Sunny, 12°C at 6pm (warmer than last week).
Crew Physical Status: Good.
EVA: 2 EVAs (n°8,9). Both went well, no issue to report.
Reports to be filed: Sol Summary, 2 EVA Requests, 2 EVA Reports, Operations Report, Journalist Report, HSO Report.
Support Requested: No support requested.